Structured Communications – Christmas 2017 Change Freeze

Structured Communications will be implementing a change freeze from Monday 18/12/2017 at 01:00 GMT until Thursday 04/01/2018 at 09:00 GMT. This freeze is implemented every year during the busy Christmas period to ensure that no changes are made to our network or infrastructure which could affect services. The freeze includes all changes to the network, … Continue reading “Structured Communications – Christmas 2017 Change Freeze”

Structured Communications will be implementing a change freeze from Monday 18/12/2017 at 01:00 GMT until Thursday 04/01/2018 at 09:00 GMT. This freeze is implemented every year during the busy Christmas period to ensure that no changes are made to our network or infrastructure which could affect services.

The freeze includes all changes to the network, non-essential maintenance, non-pre-planned change requests and infrastructure changes. Structured won’t be undertaking any planned maintenance during this time, unless we are responding to a fault.

Support tickets will not be affected by this change along with christmas changes to PBXs and current “inflight” orders already processing.

If you have any further questions regarding this freeze, please get in touch.

17-12-2017 – 22:00 till 22:10 – Planned Maintenance: Broadband

We will be upgrading our Broadband LNS gateways to the latest factory firmware release. This will require several reloads of each device as we are also upgrading the boot loaders along with the new firmware. Customer sessions are expected to drop during the above window and may flip between gateways as each one is upgraded … Continue reading “17-12-2017 – 22:00 till 22:10 – Planned Maintenance: Broadband”

We will be upgrading our Broadband LNS gateways to the latest factory firmware release. This will require several reloads of each device as we are also upgrading the boot loaders along with the new firmware.

Customer sessions are expected to drop during the above window and may flip between gateways as each one is upgraded in turn.

UPDATE 01 – 22:52
This work started late however is now complete with sessions reconnecting as expected. Anyone still without internet access are advised to power off there router / modem for 20 minutes.

LON01 – VoIP – 21/11/2017

We are aware 1 of our media gateways is not releasing channels once a call has cleared down. This is causing BUSY tones or limit exceeded messages We are currently working to resolve this ASAP. UPDATE 01 – 19:00 Emergency Works have started. Any active calls on SIP02 have been dropped. We are sorry for … Continue reading “LON01 – VoIP – 21/11/2017”

We are aware 1 of our media gateways is not releasing channels once a call has cleared down. This is causing BUSY tones or limit exceeded messages

We are currently working to resolve this ASAP.

UPDATE 01 – 19:00
Emergency Works have started. Any active calls on SIP02 have been dropped. We are sorry for any inconvenience caused.

UPDATE 02 – 19:03
SIP02 has reloaded and all services have restored. We will now look at SIP01

UPDATE 03 – 19:05
Emergency Works have started. Any active calls on SIP01 have been dropped. We are sorry for any inconvenience caused.

UPDATE 04 – 19:03
SIP01 has reloaded and all services have restored.

LON01 – C002.017 – 06/11/2017 – 11.32 – *At Risk*

Our network monitoring had alerted us to a power issues within one of our Cross-Connect / POP cabinets on the 2nd floor within Goswell Road. Our 4th floor suite is unaffected, however all services routing via C002.017 should be considered at risk until this is resolved. We are current engaged with Level3 and will provide … Continue reading “LON01 – C002.017 – 06/11/2017 – 11.32 – *At Risk*”

Our network monitoring had alerted us to a power issues within one of our Cross-Connect / POP cabinets on the 2nd floor within Goswell Road. Our 4th floor suite is unaffected, however all services routing via C002.017 should be considered at risk until this is resolved. We are current engaged with Level3 and will provide updates shortly.

UPDATE01 – 11.57
Level3 have responded to advise tier 2 Technicians are to investigate the incident and provide feedback on their findings within 1 hour.

UPDATE02 – 12:11
Level3 have responded to advise they have sent a subcase to thier techs in Goswell Road to continue to investigate.

UPDATE03 – 13:20
Level3 have issued the following statement:

A non-service affecting partial power rack failure at the Goswell Gateway in London, England is ongoing. The European NOC has advised that an electrical contractor has been engaged and is onsite working to resolve the power issue. There is no estimate as to when the power issue will be resolved; however, the European NOC has confirmed that this issue is not impacting services. Please be advised that updates for this event will be relayed at a minimum of hourly unless otherwise noted.

We do have a number of servers and core network equipment operating in redundancy mode due to the power loss and remain at risk.

UPDATE04 – 14:25
Level3 have issued the following update:

Event The European NOC has advised that the electrical contractor has replaced the faulty power cabling onsite clearing all of the alarms related to the partial power failure. At this time, efforts to contact the impacted clients to confirm service restoral are now underway.

We are still seeing network alarms in relation to power and have asked Level3 to tech to step in as we suspect one of our circuit breakers may have tripped. We will advise once further updates are given.

UPDATE05 – 18:30
We have chased Level3 for an update as no response has been given to our request. A level 1 escalation has also been raised.

UPDATE06 – 20:30
We have chased Level3 for an update as no response has been given to our request. A level 2 escalation has also been raised.

UPDATE07 – 21:58
Power has been restored and services are no longer “at risk” initial investigations advise it was more than a circuit breaker at fault. A complaint has been raised.

11/10/2017 – DSL Performance Issues

We are currently aware of an issue effecting DSL connections, we are looking into this and will update shortly. -Update: 13:24- We are starting to see connections return back to normal speeds, we are continuing to monitor and will update you shortly. -Update: 13:12- Engineers are still working on the issue. We will update with … Continue reading “11/10/2017 – DSL Performance Issues”

We are currently aware of an issue effecting DSL connections, we are looking into this and will update shortly.

-Update: 13:24-
We are starting to see connections return back to normal speeds, we are continuing to monitor and will update you shortly.

-Update: 13:12-
Engineers are still working on the issue. We will update with more information as it becomes available.

28-09-2017 – 23:00 – Planned Maintenance: Level3 – Goswell Road

We have been advised by Level3 that they will be replacing a line card between 23:00 on the 28/09/2017 & 02:00 on the 29/09/2017. This work will affect our 10GB interconnect with them as it involves the physical removal of hardware and re-termination of our fibre. Traffic will be automatically re-routed across other carriers during … Continue reading “28-09-2017 – 23:00 – Planned Maintenance: Level3 – Goswell Road”

We have been advised by Level3 that they will be replacing a line card between 23:00 on the 28/09/2017 & 02:00 on the 29/09/2017. This work will affect our 10GB interconnect with them as it involves the physical removal of hardware and re-termination of our fibre. Traffic will be automatically re-routed across other carriers during the works and limited impact is expected however traffic paths will change.

UPDATE 01
This work is complete

11-09-2017 – 21:00 – Planned Maintenance: BROADBAND via Telehouse North

We have been advised by one of our wholesale DSL providers that they will be upgrading backhaul capacity within Telehouse North on Monday the 11th September starting from 21:00 continuing until 06:00 on the 12th. We have been advised that this upgrade work may affect broadband (DSL and FTTC services) that are delivered to ourselves … Continue reading “11-09-2017 – 21:00 – Planned Maintenance: BROADBAND via Telehouse North”

We have been advised by one of our wholesale DSL providers that they will be upgrading backhaul capacity within Telehouse North on Monday the 11th September starting from 21:00 continuing until 06:00 on the 12th.

We have been advised that this upgrade work may affect broadband (DSL and FTTC services) that are delivered to ourselves via Telehouse North. DSL sessions we will be automatically re-routed via Telehouse East in the event of disruption. Users may suffer multiple PPP drops during the above window.

We apologize for any inconvenience caused

LON01 – EasyHTTP – Outbound SMTP – 14/08/2017

We are aware of outbound SMTP delivery issues on server01.r02.easyhttp.co.uk This has been caused by a compromised user account on this server that has been used to send out a large volume of SPAM this morning. The account has been terminated and is being dealt with however, it has resulted in temporary delivery restrictions from … Continue reading “LON01 – EasyHTTP – Outbound SMTP – 14/08/2017”

We are aware of outbound SMTP delivery issues on server01.r02.easyhttp.co.uk This has been caused by a compromised user account on this server that has been used to send out a large volume of SPAM this morning. The account has been terminated and is being dealt with however, it has resulted in temporary delivery restrictions from some remote hosts such as Yahoo.

We are monitoring the situation. Any failed emails depending on the remote host will be queued for re-delivery. If delivery restrictions remain then we will look to divert outbound email flow via another internal SMTP array on the sort term.

We apologize for any inconvenience caused.

UPDATE01 – 14:43
We are no longer seeing outbound emails queue on the server what have failed delivery due to the server reputation. Due to this we have agreed not to re-direct outbound SMTP. If users are still seeing bounce back emails then please email support and we will review on a case by case biases.

LON01 – Network Issue – 25/07/2017 – 14:05

We are aware of a network affecting issue. Engineers are currently working to identify the root cause and affected services. UPDATE 01 – 14:15 Following an audit of our network monitoring we can confirm this only affected our DSL services. Initial diagnostics show this to be a Zen WHOLESALE issue further upstream with one of … Continue reading “LON01 – Network Issue – 25/07/2017 – 14:05”

We are aware of a network affecting issue. Engineers are currently working to identify the root cause and affected services.

UPDATE 01 – 14:15

Following an audit of our network monitoring we can confirm this only affected our DSL services. Initial diagnostics show this to be a Zen WHOLESALE issue further upstream with one of our providers as we observed a number of L2TP Tunnels shutdown and BGP prefix withdrawals from THN.

A fault has been logged with their NOC department who confirmed that their was an issue that affected Wholesale L2TP subscribers along with there channel and direct retail customers.

Services have restored and we have provided our diagnostic information to there NOC. At this time we are unaware of the root cause and services should be considered at risk for now.

UPDATE 02 – 21:30

We are awaiting an RFO and have been advised no further repeats are expected.

LON01 – xDSL – 07/07/2017 *At Risk*

Network monitoring has advised us that one of wholesale L2TP fibre links are down for xDSL. This is part of our backup network and is not service affecting, however due to its non operation at this time the DSL network is classed as *At Risk* Further updates will be published once available. This is currently … Continue reading “LON01 – xDSL – 07/07/2017 *At Risk*”

Network monitoring has advised us that one of wholesale L2TP fibre links are down for xDSL. This is part of our backup network and is not service affecting, however due to its non operation at this time the DSL network is classed as *At Risk* Further updates will be published once available.

This is currently suspected to be a fibre break within one of our cross connects at Telehouse East.

UPDATE 01 – 07/07/2017 – 23:30
We have been advised by Telehouse staff that fibre cleaning works took place on our cross connect. This has restored service and we have confirmed this at both ends.

UPDATE 02 – 10/07/2017 – 09:30
Service has remained stable so we are closing this incident.